Archive for March, 2020

Incident: L2TP broadband

Posted: Tuesday, March 31st, 2020 at 11:46 by david labouchardiere

We are currently aware of an issue affecting authentication on L2TP broadband circuits. Engineers are already engaged and are investigating.

Further updates will be provided as and when they become available.

 

  1. david labouchardiere says:

    Engineers confirm service has now been restored. Please ask any customers still having issues to reboot their router before contacting the service desk. We apologise for any inconvenience caused.

Incident: slo24-rtr-1

Posted: Wednesday, March 18th, 2020 at 14:40 by Jordan Coleman
We are currently aware of an issue affecting the slo24-rtr-1 PE. Initial diagnostics indicate a line card failure. Engineers are investigating and engaging with our supplier. Further updates will be provided when available.
  1. Aiden Spencer says:

    Engineers have completed troubleshooting and have identified a hardware failure. New hardware is en-route to site and will be installed upon arrival, expected within the next 2 hours.

  2. Aiden Spencer says:

    The issue has been resolved and the services have now been restored.

Incident: Virgin Media Circuits

Posted: Thursday, March 5th, 2020 at 07:50 by Jonathan Clarke

We are currently seeing a number of Virgin Media circuits down. Suppliers have advised that this due to cable damage-causing wider network issues.

Suppliers have advised that engineers are at the location and fibre splicing is taking place.

Further updates will be provided on an hourly basis.

  1. Jonathan Clarke says:

    Suppliers have advised that two cables have been damaged, 96f and 160f.

    Current Situation:
    Work has commenced on the 160f to restore critical services.
    A cable gang is due onsite by 8:30 am with 2km of cable to replace the existing 96f.

    Next steps:
    New 96f cable to be pulled and spliced.

    Further updates to follow.

  2. david labouchardiere says:

    The supplier advises that work continues on the 160f to restore critical services(blue light etc)
    The original 96f still remains unavailable. We have received confirmation that the cable gang has arrived on-site and have begun their prep work to replace the existing 96f.

    Next steps
    New 96f cable to be pulled and spliced. A splicer has now arrived on-site to replace the existing splicer with further splicing resource still expected to arrive at 11am .

    Further updates to follow

  3. david labouchardiere says:

    The supplier had updated to advise that work continues on the 160f to restore services. They have seen a large number of these services come back online but there is still further work to be done on the remaining fibre tubes on this 160F Fibre.
    The original 96f still remains unavailable. The cable gang have been working on this issue all morning with splice work now beginning on the impacted 96F.

    We will continue to provide updates once available

  4. david labouchardiere says:

    The supplier advises that resplicing work is still underway with engineers continuing to work to restore services. We will continue to provide updates once they are made available

  5. david labouchardiere says:

    The supplier has advised that the fibre repair work has been completed. Any circuits still offline should try rebooting their equipment before contacting the service desk for further assistance